OpenStudy (anonymous):

My sum is -200 and the product is +6000, what two numbers would give me a sum of -200 and a product of +6000?

OpenStudy (anonymous):

x + y = -200 x * y = 6000

OpenStudy (anonymous):

can you solve that?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

how would you factor this then? 0.0025d^2-0.5d+15?

OpenStudy (anonymous):

I have no idea what that expression has to do with the problem

OpenStudy (anonymous):

you have a system of 2 equations containing 2 unknowns

OpenStudy (anonymous):

You should learn how to solve this

OpenStudy (anonymous):

do you know how to solve any system of equations? if not do you even know what it is?

OpenStudy (anonymous):

I want to factor that so I could find the roots

Directrix (directrix):

@gerba Divide each term of this: 0.0025d^2-0.5d+15 = 0 by .0025 to get d^2 - 200d + 6000 = 0 and proceed from there. I think you may have a sign error on the -.5d. c + d = -200 --> c = -200 - d cd = 6000 (-200 - d)*d = 6000 -200d - d^2 = 6000 -d^2 - 200d = 6000 d^2 + 200d = -6000 d^2 + 200d + 6000 = 0

Directrix (directrix):

Suggestion: Complete the square or use the Quadratic Formula to solve for d.